# Are you doing better than random guessing? A call for using negative controls when evaluating causal discovery algorithms

### v1

@@ 4.1 Example: Expected precision and recall for a dense 5 node DAG skeleton
$\text{E}(\text{precision}\operatorname{|}(m_{\text{max}}=10,m_{\text{true}}=8,m_{\text{est}}=7))=\frac{8}{10}=0.80$
@@
$\left(\frac{q_{(0.025,10,8,7)}}{7},\frac{q_{(0.975,10,8,7)}}{7}\right)=\left(\frac{5}{7},\frac{7}{7}\right)=(0.71,1.00).$
Hence for this DAG estimation task, it will not be highly unusual to obtain adjacency precisions as high as 1.00 under random guessing, and hence adjacency precision is not very useful for assessing performance. We can also compute the median:
$\text{median}(\text{precision})=\frac{q_{(0.5,10,8,7)}}{7}=\frac{6}{7}=0.86$
@@
Similarly, for adjacency recall we find
$\text{E}(\text{recall}\operatorname{|}(m_{\text{max}}=10,m_{\text{true}}=8,m_{\text{est}}=7))=\frac{7}{10}=0.70$
@@
$\left(\frac{k_{0.025,10,8,7}}{8},\frac{k_{0.975,10,8,7}}{8}\right)=\left(\frac{5}{8},\frac{7}{8}\right)=(0.63,0.88).$
@@
4.2 Example: F1 scores for a 5 node DAG with varying density
Figure 2: Expected F1 scores obtained under random guessing for estimating a 5 node DAG with varying true density ($m_{\text{true}}$, displayed on the horizontal axis) and varying density of the estimate ($m_{\text{est}}$, marked in color).
Figure 2 provides an overview of obtained F1 scores under random guessing across all possible combinations of estimated number of edges (marked in color) and true number of edges (horizontal axis) for 5 node DAGs. We see that it is quite possible to obtain a large F1 by random guessing if the true DAG is not very sparse, and especially, if the estimate also is not very sparse. But we stress that neither has to be overly or unrealistically dense either: For a true graph that has just 5 edges – i.e. the sparsest graph that is connected – a randomly drawn DAG with 5 edges will result in an expected F1 of 0.5, and placing all 10 possible edges results in an F1 score of 0.66. If we instead consider a more dense graph, e.g. a true DAG with 8 edges, we are back in then scenario already considered above, and we see that we can find a peak F1 score of 0.89 by placing all edges.
@@ 5 A test of overall skeleton fit
This is done by comparing the observed number of true positives, $TP_{obs}$ with the appropriate hypergeometric distribution. Formally, if we let $X\sim\text{HyperGeom}(m_{\text{max}},m_{\text{true}},m_{\text{est}})$, a one-sided p-value for $H_{0}$ is computed as
$P(X\geq TP_{obs})$
i.e., the probability of getting as least as many true positives as the observed number, if edges were in fact randomly placed.
Note than since the text is exact (and based on a discrete probability distribution), it will be conservative.
5.1 Application on Metropolit cohort dataset
@@
Table 3: Adjacency confusion matrix replicated from Petersen et al. (2023a).
We will reanalyze data from Petersen et al. (2023a). Here, the temporal PC algorithm (TPC) was used on a cohort data set of $n=3145$ Danish men to identify possible causes of depression and heart disease, as well as their interplay. Two experts were also asked to construct a model for the data based on existing studies and subject-field knowledge, and their DAG was compared to the output of TPC. For the comparison here, we assume that the expert model is correct and wish to evaluate if TPC performs better than a random control at estimating the expert model. The DAGs have 22 nodes, and hence $m_{\text{max}}=231$ possible edges.
@@
We conduct an overall test of skeleton fit by comparing the obtained number of true positives, $\text{TP}_{\text{obs}}=10$, with $\text{HyperGeom}(231,30,30)$ and we find $p=0.002$. Hence, we reject $H_{0}$ and conclude that TPC performs significantly better than random guessing in this application.
@@ 6 Simulation-based negative controls for more general metrics
One example is conditional orientation metrics (also sometimes referred to as arrowhead metrics) (see e.g. Andrews et al. (2019)). These metrics describe correct orientation among correctly placed edges. We conjecture that simple exact distributional results under random guessing do not exist for this classification task. The main issue is that placement of oriented edges does not correspond to an urn experiment analogue, as consecutive placement steps are not independent when the goal is to output e.g. a valid DAG: If we have already placed oriented edges such that $X\to Y\to Z$, it is (no longer) possible to have an edge pointing from $Z$ to $X$, as this would introduce a cycle and the graph would thus no longer be a valid DAG. Thus, describing expected behavior under random guessing when also taking edge orientations into account is more complicated.
However, we can easily use simulation to obtain an empirical estimate of the distribution of a given metric under random edge placement – oriented or not. We propose the following procedure:
@@ 1. Standard simulation study:
Conduct the simulation study as usual (focusing on one or more metrics of interest) and for each simulation repetition, store the true (simulated) DAG as well as the density (or number of edges) for the estimated DAG.
2. Negative control estimation:
Draw a large (e.g. 1000) number of random DAGs with densities sampled from the distribution obtained in Step 1.
@@ 3. Negative control evaluation:
Compare each of the random DAGs to the true DAGs from step 1 by computing the metric(s) of interest. Report the mean(s) as the expected performance under random guessing, and use the empirical quantiles to construct e.g. a 95% confidence interval.
@@ 4. Comparison:
Finally, compare the metrics obtained under random guessing with what is obtained from the algorithm being evaluated. In order to draw statistical inference, compute their pairwise performance differences for each ”true” DAG, and construct e.g. a 95% confidence interval using the empirical quantiles of the distribution of these differences.
We here focus on the case of a simulation study, where many different ground truths are simulated, but in Section 6.2 we also provide an example of how to adapt the procedure to be suited for evaluation of a real data application where there is only a single ground truth.
@@
Note that it is important for obtaining valid statistical inference that it is conducted on the pair-wise differences in performance, as inference e.g. based on whether or not confidence intervals overlap is highly conservative (Knol et al., 2011).
6.1 Example: Structural Hamming distance in simulation study
The structural Hamming (SHD) distance measures the number of edge reversals, removals and additions in order to transform one graph (estimate) into another (truth). This metric hence includes both adjacency information and orientation information, and thus the results from Section 3 do not apply. We will instead compute an SHD negative control for evaluating the PC algorithm by use of simulation. We choose a simulation scenario, where it is well-known that PC struggles, namely a dense true data generating mechanism, and hence we do not expect PC to do better than a random control.
We proceed as follows (following the steps from above):
1. Standard simulation study:
We draw 1000 random graphs over 10 nodes, each with 30 edges (i.e., quite dense). From each DAG, we simulate 400 independent Gaussian observations with randomly drawn regression parameters and error variances [^1] We use default options for regression parameters and error variances from the simGausFromDAG() function in the causalDisco R package (Petersen, 2022).. For each dataset, we apply the PC algorithm with test significance level $\alpha=0.05$.
  1. (a)
    The metric of interest, SHD, is computed for each of the 1000 graphs, and we find a mean SHD of $27.3$ with a 95% confidence interval of $(21;33)$ (based on the empirical quantiles).
  2. (b)
    We store the ”true” DAGs as well as the distribution of number of edges across the 1000 applications of the PC algorithm. The number of edges range from 7 to 19 with a mean of $13.4$.
2. Negative control estimation:
We draw 1000 random DAGs over 10 nodes with number of edges independently sampled from the distribution from Step 1 (b).
3. Negative control evaluation:
For each of the 1000 negative controls, we compare with the ”true” DAGs from Step 1 by computing the SHD. We find a mean SHD of $30.1$ (95% CI: $24;35$) for the negative controls.
4. Comparison:
We find the mean (pairwise) difference in SHD for PC vs. random controls to be $2.8$, 95% CI: $(-5;11)$. As 00 is included in this confidence interval, we conclude that PC is not performing statistically significantly different from a random control in this scenario (at a 5% test level).
6.2 Application: Negative control for evaluation of NOTEARS on Sachs data
In this example, we evaluate a causal discovery procedure using a (single) real data set with a known ground truth. Hence, we only have a single ”true” DAG (and a single value of $m_{\text{est}}$), but we will again use repeated negative control simulations in order to gain more stable results and allow for computing a confidence interval.
@@ 6.2 Application: Negative control for evaluation of NOTEARS on Sachs data
We consider evaluation of the NOTEARS algorithm on the Sachs dataset (Sachs et al., 2005), which is often used to evaluate causal discovery algorithms. The ground truth DAG for the Sachs dataset has 11 nodes and 20 edges (although there also exists an alternative version with only 17 edges) [^2] We use this ”truth” version of the Sachs dataset because that is also what was used to evaluate the NOTEARS algorithm. We obtain the ground truth from the repository https://github.com/cmu-phil/example-causal-datasets.
In the original article proposing NOTEARS, the authors report that NOTEARS obtains an SHD of 22 on the Sachs dataset by placing 16 edges (Zheng et al., 2018). To compute performance on a negative control for this task, we simulate 1000 random DAGs over 11 nodes with 16 edges (i.e., same number of edges as NOTEARS). The average SHD over the random controls, comparing to the Sachs ground truth, is $27.4$ with a 95% confidence interval of $(22,33)$. Hence, testing at a 5% level, NOTEARS does not obtain significantly better SHD performance than the random control on the Sachs dataset (although the value lies on the border of the confidence interval).
@@ 7 Discussion
The distributional results for adjacency metrics presented in Section 3 are conditional on three quantities: The maximal number of edges in the DAG ($m_{\text{max}}$), the number of edges in the true DAG ($m_{\text{true}}$) and the number of edges in the estimated DAG ($m_{\text{est}}$). Clearly, conditioning on the first two is completely uncontroversial: $m_{\text{max}}$ is a mathematical property of the graph, and in order to do causal discovery, we have to assume that there exists an (unknown) true DAG with a set number of edges ($m_{\text{true}}$). But conditioning on $m_{\text{est}}$ may be debated. Our motivation for doing so is as follows: Many causal discovery algorithms – e.g. PC (Spirtes and Glymour, 1991), FCI (Spirtes et al., 2000), GES (Chickering, 2002), and GRaSP (Lam et al., 2022) – require choosing a tuning parameter, which will in practice directly control the number of outputted edges (e.g. test significance level in constraint-based algorithms or score penalties in score-based algorithms). The way causal discovery algorithms are mostly applied, this tuning parameter is not chosen in a data-driven manner, but rather set at somewhat arbitrary ”standard” values. Alternatively, in some instances, it may be chosen based on external background knowledge (Petersen et al., 2023a). In either case, the number of edges in the estimated graph is de facto chosen a priori, in which case we can meaningfully condition on it.
@@
Some work has been proposed for data-driven tuning of causal discovery algorithms (Biza et al., 2020). If such methods are applied, the density of the estimated graph will generally be estimated from data. In this case, we lose the distributional results for the adjacency metrics presented in Section 3, as the number of true positives will no longer be hypergeometrically distributed. But we can still use the proposed simulation-based pipeline from Section 6 to obtain a negative control for such algorithms as well.
@@
Another reason for preferring to condition on $m_{\text{est}}$ is related to interpretability. By only considering one value of $m_{\text{est}}$, we compare our causal discovery procedure with a well-specified negative control experiment, namely placing the same number of edges at random. This argument also has implications for how we ought to compare two different causal discovery algorithms, namely, by tuning them to estimate the same number of edges and then compare their outputs. Otherwise, we may be comparing sparse outputs with dense ones, and as we have seen above, such a comparison may not be meaningful, and will definitely be difficult to interpret. Moreover, by tuning them to produce equally dense outputs, we also remove tuning parameters from the evaluation equation altogether, and replace this parameter with the simpler and more interpretable notion of outputted graph density.
@@
However, we want to stress that it can be problematic to consider only a single density in a simulation-based evaluation study if the algorithm being evaluated is able to thereby learn the (unique) intended density (Petersen et al., 2023b). This evaluation design flaw has been present for several supervised discovery algorithms (Li et al., 2020; Xu and Xu, 2021; Yu et al., 2019), and could harm transportability greatly. We propose that a range of densities should hence always be considered when conducting simulation-based evaluation studies of discovery algorithms that may learn the density directly from training data. But to ease interpretation, the results should ideally be presented stratified according to density.
@@
The work presented here has focused on Erdös-Rényi type graphs. This assumption is important for the distributional results in Sections 3-5, as the hypergeometric distribution requires random draws. Non-central versions of the hypergeometric distribution allows for biased draws of edges, but we do not believe this is very useful for describing causal graphs: It would allow certain edges to be more likely to be present than others, but would still not consider graph properties beyond singular edges and hence not be appropriate for describing e.g. graphs that exhibit clustering. However, if a specific evaluation study wants to target such graphs, the simulation-based method proposed in Section 6 can straightforwardly be applied, simply by simulating random graphs from the intended target graph type.
@@
As mentioned in Section 5, the exact distributional results for adjacency metrics will by definition result in conservative statistical inference, i.e. conservative control of type I error in statistical tests and overly wide confidence intervals. Due to the discrete nature of the hypergeometric distribution, this is especially pronounced when $m_{\text{max}}$ is small, i.e. when there are only few nodes. However, we argue that the considered null hypothesis is very crude – assuming completely random replacement of $m_{\text{est}}$ edges – and hence we do not consider conservative inference to be very problematic. Informally, we would ideally like to perform markedly better than random guessing, not just borderline significantly so!

### v2

@@ 4.1 Example: Expected Adjacency Precision and Recall for a Dense 5 Node DAG
$\text{E}(\text{prec}\operatorname{|}m_{\text{max}}=10,m_{\text{true}}=8,m_{\text{est}}=7)=\frac{8}{10}=0.80$
@@
$\left(\frac{q_{(0.025,10,8,7)}}{7},\frac{q_{(0.975,10,8,7)}}{7}\right)=\left(\frac{5}{7},\frac{7}{7}\right)\simeq(0.71,1.00).$
Hence for this DAG estimation task, it will not be highly unusual to obtain adjacency precisions as high as 1.00 under random guessing, and thus adjacency precision is not very useful for assessing performance. We can also compute the median precision:
$\text{median}(\text{prec})=\frac{q_{(0.5,10,8,7)}}{7}=\frac{6}{7}\simeq 0.86$
@@
For adjacency recall we find
$\text{E}(\text{recall}\operatorname{|}m_{\text{max}}=10,m_{\text{true}}=8,m_{\text{est}}=7)=\frac{7}{10}=0.70$
@@
$\left(\frac{q_{0.025,10,8,7}}{8},\frac{q_{0.975,10,8,7}}{8}\right)=\left(\frac{5}{8},\frac{7}{8}\right)\simeq(0.63,0.88).$
@@
4.2 Example: Adjacency F1 Scores for a 5 Node DAG With Varying Density
Figure 2: Expected Adjacency F1 Scores Under Random Guessing for Estimating 5 Node DAGs.
Figure 2 provides an overview of obtained F1 scores under random guessing across all possible combinations of estimated number of edges (horizontal axis) and true number of edges (marked in color) for 5 node DAGs. We see that it is quite possible to obtain a large F1 score by random guessing if the true DAG is not very sparse, and especially, if the estimate is also not very sparse. But we stress that neither has to be overly or unrealistically dense either: For a true graph that has just 5 edges — i.e., only one edge more than the sparsest graph that is connected — a randomly drawn DAG with 5 edges will result in an expected F1 of 0.5, and placing all 10 possible edges results in an F1 score of 0.66. If we instead consider a more dense graph, e.g., a true DAG with 8 edges, we are back in the scenario already considered above, and we see that we can find a peak F1 score of 0.89 by placing all edges.
@@ 5 A Test of Overall Skeleton Fit
This is done by comparing the observed number of true positives, $\textit{tp}_{\text{obs}}$ with the appropriate hypergeometric distribution. Formally, if we let $X\sim\text{HyperGeom}(m_{\text{max}},m_{\text{true}},m_{\text{est}})$, a one-sided p-value for $H_{0}$ is computed as
$P(X\geq\textit{tp}_{\text{obs}})$
i.e., the probability of getting at least as many true positives as the observed number if edges were in fact randomly placed. Note that since the test is exact (and based on a discrete probability distribution), it will be conservative.
5.1 Application: Temporal PC on the Metropolit Cohort Dataset
Table 4: Adjacency Confusion Matrix Replicated From Petersen et al. [2023a].
@@ 5.1 Application: Temporal PC on the Metropolit Cohort Dataset
We will reanalyze data from Petersen et al. [2023a]. In that study, the temporal PC algorithm (TPC) [Petersen et al., 2021] was used on a cohort data set of $n=3145$ Danish men to identify possible causes of depression and heart disease, as well as their interplay. Two experts were also asked to construct a model for the data based on existing studies and subject-field knowledge, and their DAG was compared to the output of TPC. For the comparison here, we assume that the expert model is correct and wish to evaluate if TPC performs better than a negative control at estimating the expert model. The DAGs have 22 nodes, and hence $m_{\text{max}}=231$ possible edges.
@@
We conduct an overall test of skeleton fit by comparing the obtained number of true positives, $\textit{tp}_{\text{obs}}=10$, with $\text{HyperGeom}(231,30,30)$ and we find $p=0.002$. Hence, we reject $H_{0}$ and conclude that TPC performs significantly better than random guessing in this application.
@@ 6 Simulation-Based Negative Controls for More General Metrics
One example is conditional orientation metrics (also sometimes referred to as arrowhead metrics) (see e.g., Andrews et al. [2019]). These metrics describe correct orientations among correctly placed edges. We conjecture that simple exact distributional results under random guessing do not exist for this classification task. The main issue is that consecutive edge placement steps are not independent when the goal is to output e.g., a valid DAG: If we have already placed oriented edges such that $X\to Y\to Z$, it is no longer possible to have an edge pointing from $Z$ to $X$, as this would introduce a cycle and the graph would then no longer be a valid DAG. Thus, describing expected behavior under random guessing when also taking edge orientations into account is more complicated.
However, we can easily use simulation to obtain an empirical estimate of the distribution of a given metric under random edge placement — oriented or not. Let $b$ be the number of repetitions in the simulation study, and let $f$ denote some metric of interest.We propose the following procedure:
@@ 1. Standard simulation study:
Conduct the simulation study as usual: Simulate $b$true" DAGs $G_{\text{true}}^{1},...,G_{\text{true}}^{b}$, generate appropriate data for each, and use the causal discovery algorithm of interest to obtain estimated graphs $\hat{G}_{\text{algo}}^{1},...,\hat{G}_{\text{algo}}^{b}$. For each $i\in\{1,...,b\}$, compare the true and estimated graphs by computing the metric of interest, $f(G_{\text{true}}^{i},\hat{G}_{\text{algo}}^{i})$, and for each estimated graph $\hat{G}_{\text{algo}}^{i}$, count the number edges, $m_{\text{est}}^{i}$.
2. Negative control simulation:
For each $i\in\{1,...,b\}$, draw a negative control random DAG $\hat{G}_{\text{NC}}^{i}$ with number of edges sampled randomly from $\{m_{\text{est}}^{1},...,m_{\text{est}}^{b}\}$ (with replacement).
@@ 3. Negative control evaluation:
Compare each negative control with the corresponding true graph by computing the metric of interest, $f(G_{\text{true}}^{i},\hat{G}_{\text{NC}}^{i})$. Report the mean as the expected performance under random guessing, and use the empirical quantiles to construct e.g., a 95% confidence interval.
@@ 4. Comparison:
Finally, compare the metrics obtained under random guessing with the metrics obtained for the evaluated algorithm. In order to draw statistical inference, consider pairwise comparisons and conduct a one-sided statistical test. Compute the $p$-value as
$p=\frac{1}{b}\sum_{i=1}^{b}\mathbf{1}\left(f(G_{\text{true}}^{i},\hat{G}_{\text{algo}}^{i})\leq f(G_{\text{true}}^{i},\hat{G}_{\text{NC}}^{i})\right)$
for metrics where small values are favorable (otherwise reverse the inequality).
@@
Note that it is important for obtaining valid statistical inference that it is conducted on the pairwise comparisons of performance, as inference e.g., based on whether or not confidence intervals overlap is highly conservative [Knol et al., 2011].
If the evaluated algorithm does not estimate a DAG, we suggest that Step 2 is altered to match the output of the evaluated algorithm. For example, if the algorithm only aims to learn the Markov equivalence class of the data generating DAG, as represented by a CPDAG, we would simply use negative control CPDAGs in Step 2 by first drawing DAGs and subsequently finding their encompassing CPDAGs.
We have here focused on the case of a simulation study where many different ground truth graphs are simulated in Step 1, but in Section 6.2 we also provide an example of how to adapt the procedure to be suited for evaluation of a real data application where there is only a single ground truth.
6.1 Example: Simulation Study Evaluating the PC Algorithm
Table 5: Comparisons of PC Algorithm and Negative Controls.
PC Negative control
Mean CI Mean CI $p$
Dense case ($m_{\text{true}}=30$)
   SHD 27.33 $(21,33)$ 31.23 $(26,36)$ 0.202
   Adjacency precision 0.85 $(0.65,1.00)$ 0.66 $(0.42,0.87)$ 0.122
   Adjacency recall 0.38 $(0.27,0.50)$ 0.29 $(0.17,0.43)$ 0.245
   Orientation precision 0.65 $(0,1)$ 0.50 $(0,1)$ 0.360
   Orientation recall 0.40 $(0.00,0.78)$ 0.37 $(0.00,0.78)$ 0.464
   Proportion recovered v-structures 0.05 $(0.0,0.2)$ 0.02 $(0.00,0.14)$ 0.563
   SID (lower bound) 67.73 $(46,83)$ 74.23 $(56,85)$ 0.317
   SID (upper bound) 79.48 $(61,90)$ 79.10 $(63,88)$ 0.557
Sparse case ($m_{\text{true}}=15$)
   SHD 10.1 $(4,15)$ 21.30 $(17,25)$ 0.002
   Adjacency precision 0.9 $(0.73,1.00)$ 0.33 $(0.091,0.571)$ 0.000
   Adjacency recall 0.7 $(0.47,0.87)$ 0.25 $(0.067,0.467)$ 0.001
   Orientation precision 0.9 $(0.5,1.0)$ 0.52 $(0,1)$ 0.273
   Orientation recall 0.5 $(0.00,0.91)$ 0.36 $(0,1)$ 0.316
   Proportion recovered v-structures 0.3 $(0.0,0.8)$ 0.01 $(0.00,0.14)$ 0.106
   SID (lower bound) 29.3 $(7,55)$ 51.01 $(29,74)$ 0.072
   SID (upper bound) 51.5 $(22,81)$ 58.43 $(36,81)$ 0.350
[…]
We showcase the proposed simulation-based negative control procedure by evaluating the performance of the PC algorithm [Spirtes and Glymour, 1991]. We construct a small toy simulation study considering the task of learning 10-node DAGs (or more specifically, CPDAGs corresponding to their Markov equivalence classes) from linear Gaussian data generated according to the DAGs. This is a scenario where the PC algorithm is sound and complete in the large sample limit, so in principle we should expect good performance. However, it is well-known that PC struggles on finite data when the true data generating mechanism is dense, because the algorithm is biased towards sparse graphs (see e.g., [Petersen et al., 2023b]). We therefore consider a moderate sample size of $n=400$ observations, and two settings for how dense the true DAGs are: A dense case with $m_{\text{true}}=30$ edges and a sparse case with $m_{\text{true}}=15$. We expect that PC performs better than negative controls in the sparse case, but not in the dense case.
To evaluate PC, we consider five different metrics that make use of orientation information (and are hence beyond the scope of Sections 3-5): Structural Hamming distance, orientation precision, orientation recall, proportion recovered v-structures, and structural intervention distance (SID) lower and upper bounds [^2] The SID can only be reported as bounds as we are comparing a true DAG with an estimated CPDAG, see details in Supplementary Materials C.. For comparability with the previous sections, we also report two adjacency metrics considered above; namely adjacency precision and recall. We provide definitions of the metrics and additional details about the simulation study in Supplementary Materials C.
Table 5 presents the results. In the dense case, we find that none of the metrics have sufficient discriminatory power to distinguish between PC and the negative control (testing at e.g., a 5% significance level). For this case, PC estimated graphs with numbers of edges ranging from 7 to 19 with a mean of $13.3$, which is severely biased towards sparsity, as expected.
For the sparse case, the number of edges estimated by PC ranges from 7 to 16 with a mean of $11.5$, i.e., a better match with $m_{\text{true}}$. We find that some metrics show significant differences between PC and the negative control, while others do not: SHD, and adjacency precision and recall are significantly better for PC than the negative controls (at a 5% level), while the others are not. While the mean values for PC and negative controls are generally quite far apart, the 95% confidence intervals reveal very broad distributions for orientation precision, orientation recall, proportion recovered v-structures and both SID bounds. Hence, there is a lot of variability in these metrics — both for PC and negative controls — and they are thus perhaps not very useful for evaluating this case. Another takeaway from this application is that reporting only mean values of metrics is not advisable; ideally, some description of the distribution (e.g., confidence intervals) should be included.
Overall, we find that including negative controls can thus also be used to provide insights into the level of informativeness of a specific metric in scenarios where we have an established consensus of whether a certain causal discovery procedure "works well" or not.
6.2 Application: Structural Hamming Distances on the Sachs Data
Table 6: Structural Hamming Distances for the Sachs Data.
Observed Negative control
Algorithm SHD $m_{\text{est}}$ Mean SHD $p$
PC 23 24 31.54 0.001
NOTEARS 22 16 27.10 0.050
LiNGAM 30 33 34.43 0.083
GES 30 30 34.24 0.114
BOSS 35 32 35.24 0.510
@@ 6.2 Application: Structural Hamming Distances on the Sachs Data
In this application, we consider the Sachs dataset [Sachs et al., 2005], which is often used to evaluate causal discovery algorithms. The ground truth DAG for the Sachs dataset has 11 nodes and $m_{\text{true}}=20$ edges [^3] Note that there also exists an alternative version with only 17 edges..
We evaluate the performance of five causal discovery procedures, namely PC [Spirtes and Glymour, 1991], GES [Chickering, 2002], LiNGAM [Shimizu et al., 2006], NOTEARS [Zheng et al., 2018] and BOSS [Andrews et al., 2023]. We apply each of these algorithms to the Sachs dataset and compute their SHD and estimated number of edges, $m_{\text{est}}$. Based on $m_{\text{est}}$, we simulate 1000 negative controls separately for each algorithm, and report the mean SHD over the negative controls. For algorithms that return a DAG (LiNGAM and NOTEARS), we simulate negative control DAGs, and for algorithms that return a CPDAG (PC, GES, and BOSS), we simulate negative control CPDAGs. In all cases we compare to the ground truth DAG by computing a one-sided $p$-value testing how often the discovery algortihm performs at least as well as the negative control (according to SHD). Additional details about the application are provided in Supplementary Materials D.
Table 6 summarizes the results. We see that while the smallest SHD value is obtained by NOTEARS, PC produces the smallest $p$-value ($p=0.001$) and hence is the furthest removed from random guessing, seconded by NOTEARS ($p=0.050$) and LiNGAM ($p=0.083$). In the other end of the spectrum, we find that GES, and especially BOSS, are not significantly different from random guessing testing at e.g., a 10% significance level.
This application illustrates that it is not very meaningful to compare and interpret differences in SHD without taking into account the number of edges placed. A lower SHD does not necessarily mean that an algorithm is further removed from random guessing; it may just reflect a more preferable level of sparsity in the estimated graph, which can be an artefact of the chosen metric. For example for SHD on the Sachs dataset, we can obtain $\text{SHD}=m_{\text{true}}=20$ — i.e., a value that outperforms all considered algorithms — simply by "estimating" the empty graph. Hence, a more meaningful ranking may come about by considering the size of the negative control $p$-values.
@@ 7 Discussion
The results presented here were developed with the aim of evaluating algorithms in an artificial "lab" setting where we have access to a known ground truth. Subsequent to such evaluations, the algorithms should of course also be tested in practice in real data applications without a known ground truth graph. How to asses performance in such scenarios is fundamentally difficult, as causal discovery is a unsupervised problem. One approach for validating causal discovery on real world data is to compare graphs (or resulting effect estimates) found by causal discovery algorithms with expert-made graphs based on theory or existing literature [Petersen et al., 2023a, Gururaghavendran and Murray, 2024]. However, such comparisons are of course only feasible when a comprehensive body of knowledge about the considered variables already exists, and even in this case, it is highly time consuming to construct expert graphs. A more broadly applicable evaluation strategy has been proposed by Eulig et al. [2024]. They compare conditional independence fit of a candidate DAG (given by experts or estimated using causal discovery) with a baseline obtained by randomly permuting nodes. By doing so, they construct a statistical test of fit with a similar interpretation to the test proposed in Section 5, although focusing on conditional independencies implied by the DAG rather than edge presence directly. However, as the authors note, such a test is not readily applicable for validating causal discovery fit, as most algorithms use conditional independence information for estimating the graph, and hence a subsequent test based on the same information would result in overfitting. This could however be resolved by applying data splitting if the sample size renders such an approach feasible.
@@
The distributional results for adjacency metrics presented in Section 3 are conditional on three quantities: The maximal number of edges in the DAG ($m_{\text{max}}$), the number of edges in the true DAG ($m_{\text{true}}$) and the number of edges in the estimated DAG ($m_{\text{est}}$). Clearly, conditioning on the first two is completely uncontroversial, but conditioning on $m_{\text{est}}$ may be debated. Our motivation for doing so is as follows: Many causal discovery algorithms — e.g., PC [Spirtes and Glymour, 1991], FCI [Spirtes et al., 2000], GES [Chickering, 2002], and GRaSP [Lam et al., 2022] — require choosing a tuning parameter, which will in practice directly control the number of outputted edges (e.g., test significance level in constraint-based algorithms or score penalties in score-based algorithms). Although we do not generally have a characterization of the relationship between the tuning parameter values and the resulting value of $m_{\text{est}}$, the relationship is generally deterministic in a single causal discovery application. The way causal discovery algorithms are mostly applied, the tuning parameter is not chosen in a data-driven manner, but rather set at somewhat arbitrary "standard" values. Alternatively, in some instances, it may be chosen based on external background knowledge [Petersen et al., 2023a]. In either case, the number of edges in the estimated graph is de facto chosen a priori, in which case we can meaningfully condition on it.
@@
Some work has been proposed for data-driven tuning of causal discovery algorithms [Biza et al., 2020]. If such methods are applied, $m_{\text{est}}$ will generally be estimated from data. In this case, we lose the distributional results for the adjacency metrics presented in Section 3, as the number of true positives will no longer be hypergeometrically distributed. But we can still use the proposed simulation-based pipeline from Section 6 to obtain a negative control for such algorithms.
@@
Another reason for preferring to condition on $m_{\text{est}}$ is related to interpretability. By only considering one value of $m_{\text{est}}$, we compare our causal discovery procedure with a well-specified negative control condition, namely placing the same number of edges at random. This argument also has implications for how we ought to compare two different causal discovery algorithms; to increase interpretability, we advise to either tune the algorithms to estimate the same number of edges and then compare their outputs, or follow the strategy from Section 6.2 and compare negative control $p$-values. Otherwise, we may be comparing sparse outputs with dense ones without accounting for the different difficulties in estimating sparse and dense graphs, and as we have seen above, such a comparison may not be meaningful, and will definitely be difficult to interpret. Tuning algorithms to produce equally dense outputs has an additional benefit by removing the (difficult to interpret) tuning parameters from the evaluation equation altogether, and replacing them with the simpler notion of outputted graph density.
@@
However, we want to stress that it can be problematic to consider only a single density in a simulation-based evaluation study if the algorithm being evaluated is able to thereby learn the (unique) intended density [Petersen et al., 2023b]. This evaluation design flaw has been present for several supervised discovery algorithms [Li et al., 2020, Xu and Xu, 2021, Yu et al., 2019], and could harm transportability greatly. We propose that a range of densities should therefore always be considered when conducting simulation-based evaluation studies of discovery algorithms that may learn the density directly from training data. But to ease interpretation, the results should ideally be presented stratified according to density.
We have considered a range of different metrics that may be used to evaluate causal discovery methods, but the list is clearly not exhaustive. We have focused on edgewise and structural evaluations, as these are most commonly reported [Gentzel et al., 2019], but we advice that all evaluation studies should include a critical consideration of what metrics are relevant for the specific intended use case. For example, if an intended usecase is mostly focused on causal information flow, and not whether effects are direct or indirect, it is natural to consider a metric that explicitly counts preserved ancestral information [Bang et al., 2024]. Or, if one is interested in using a causal discovery estimate for subsequent effect estimation and inference, one should include metrics on the intervention distribution, possibly targeting a specific causal estimand of interest [Gentzel et al., 2019].
@@
The work presented here has focused on Erdős-Rényi type graphs. This assumption is important for the distributional results in Sections 3 - 5, as the hypergeometric distribution requires random draws. Non-central versions of the hypergeometric distribution allows for biased draws of edges, but we do not believe this is very useful for describing causal graphs: It would allow certain edges to be more likely to be present than others, but would still not consider graph properties beyond singular edges and hence not be appropriate for describing for example graphs that exhibit clustering. However, if a specific evaluation study wants to target such graphs, the simulation-based method proposed in Section 6 can straightforwardly be applied, simply by simulating random graphs from the intended target graph type, see e.g., [Albieri and Didelez, 2014].
@@
As mentioned in Section 5, the exact distributional results for adjacency metrics will by definition result in conservative statistical inference, i.e. conservative control of type I error in statistical tests and overly wide confidence intervals. Due to the discrete nature of the hypergeometric distribution, this is especially pronounced when $m_{\text{max}}$ is small, i.e. when there are only few nodes. However, we argue that the considered null hypothesis is very crude — assuming completely random replacement of $m_{\text{est}}$ edges — and hence we do not consider conservative inference to be very problematic. Informally, we would ideally like to perform markedly better than random guessing, not just borderline significantly so!